The average daily sales of 700 branch office was rs 120 thousands and the standard deviation is rs 15,000 during the distribution to be normal how many branch have sale


1
Expert's answer
2021-06-07T13:28:31-0400

Assuming the distribution to be normal indicate how many branches have sales between 1. Rs. 90 thousands and Rs 115 thoudands

Given "\mu=120 \ thousands, \sigma=15 \ thousands"


"P(90<X<115)=P(X<115)-P(X\leq 90)"

"=P(Z<\dfrac{115-120}{15})-P(Z\leq \dfrac{90-120}{15})"

"\approx P(Z<-0.3333)-P(Z\leq -2)"

"\approx0.36944-0.02275\approx0.3467"

"0.3467(700)=243"

243 branches have sales between "Rs.90,000" and "Rs.115,000."


2. Rs. 110 thousands and Rs 135 thoudands

Given "\mu=120 \ thousands, \sigma=15 \ thousands"


"P(110<X<135)=P(X<135)-P(X\leq 110)"

"=P(Z<\dfrac{135-120}{15})-P(Z\leq \dfrac{110-120}{15})"

"\approx P(Z<1)-P(Z\leq -0.6667)"

"\approx0.84134-0.25249\approx0.5889"

"0.5889(700)=412"

412 branches have sales between "Rs.110,000" and "Rs.135,000."